# Driving of an open quantum system at finite temperature across first- and second-order quantum phase transitions

### v2

@@ 3.3 Target states and fidelity
The target state for any of the above-described driving procedures is the ground state of the qubit system at the final parameter point ${\mathbb{\Lambda}=\mathbb{\Lambda}_{\rm F}}$. For the path across the first-order QPT with ${\mathbb{\Lambda}_{\rm F}=(0.25,1.2)}$, the final ground state is given by a single eigenvector $|E_{0}(\mathbb{\Lambda}_{\rm F})\rangle$, which is well energetically separated from the other eigenvectors. So, if $\hat{\rho}_{\text{S}}(t_{\rm F})$ is the density operator of the qubit system at the end of the driving procedure, see Eq. (15), then the fidelity of the target-state preparation is
@@
$\mathcal{F}_{2}={\rm tr}\left[\hat{P}_{{\rm gs}}(\mathbb{\Lambda}_{\rm F})\,\hat{\rho}_{\text{S}}(t_{\rm F})\right]=$
@@ 3.4 Notes on the HEOM calculations
The evolution of the qubit density operator $\hat{\rho}_{\text{S}}(t)$ from Eqs. (14) and (15) is determined by the HEOM method Tani89 ; Tani06 ; Fruch16 ; Tani20 ; Lamb23 , which is, in principle, capable to yield exact description of the system dynamics. In particular, it goes beyond the approximations involved in the common Lindblad formalism, which strictly relies on the assumption that environment-induced processes on the system are Markovian. The HEOM method leads to an extensive set of first-order coupled differential equations to be solved. We do not intend here to derive these equations or to immerse deeper into the description of the method, we just introduce the parameters involved in the forthcoming calculations.
In short, the HEOM method requires the bath correlation function (2.2) to be expanded as a complex series of exponentials. For the Drude-Lorentz distribution (11) the series takes the form
@@
with coefficients ${c_{0}\in{\mathbb{C}}}$ and ${c_{1},c_{2},\ldots\in{\mathbb{R}}}$ depending on parameters $q$ and $\gamma$, and on the temperature $T$. The explicit expressions are given, e.g., in Ref. Lamb23 . In numerical calculations the sum in Eq. (26) can be truncated at its $M$th term. For low temperatures, the convergence is slow and $M$ must be relatively large to reach an acceptable precision of calculations. On the other hand, for high temperatures even very small $M$ may lead to reasonable results. In any case, to compensate this truncation, a so-called terminator, i.e, a term that speeds up the convergence, is included.
The cornerstone of HEOM is the use of auxiliary density matrices, from which the desired system density operator $\hat{\rho}_{\text{S}}(t)$ is determined. These auxiliary matrices are not physical but represent just mathematical tools introduced for computational purposes to take into account non-Markovian effects. The amount of these matrices depends on the number of layers $L$ chosen accordingly to the coupling strength $q$ and time scale $\tau_{\rm E}$ to guarantee the convergence.
@@ 4 RESULTS
In our calculations we use the value ${\gamma=10}$ for the width of the Drude-Lorentz distribution (11) in units of $\varepsilon$. This means that typical time scale of the environment is smaller (10 times for ${\lambda=\chi=0}$) than that of the qubit system. We consider three values (again in units of $\varepsilon$) of the effective system-environment coupling strengths: ${q=0}$ (no coupling), ${q=0.01}$ (weak coupling) and ${q=0.1}$ (strong coupling). In most calculations, we set the size of the qubit system to a relatively low value, namely ${N=10}$. In that case, the width $\gamma$ of the Drude-Lorentz distribution covers the whole spectrum of the qubit system with ${\lambda=\chi=0}$.
When $q\neq 0$, at low temperature, the bath correlation function $C(t)$ becomes negative for a short lapse of time, inducing a non-Markovian behavior. Therefore, values up to ${M=12}$ for the cutoff parameter and ${L=3}$ for the number of layers are needed to assure the convergence of the HEOM method. In the high-temperature case this is not longer true and ${M=4}$ and ${L=2}$ are enough to assure the convergence.
@@
where $(N+1)$ is the dimension of the qubit Hilbert space. Therefore, we consider the fidelity for different driving times $t_{\rm F}$ as a function of temperature $T$ for ${T\leq 25}$.
@@ 4.1 First-order QPT
Figure 3: Final fidelity of the ${N=10}$ qubit system for drivings across the first-order QPT precursor. Results for driving protocols A and B are shown in the top and bottom rows, respectively. The effective coupling $q$ and the interaction operator ${\hat{Q}=\hat{J}_{x},\hat{J}_{z}}$ are indicated above or within each subplot. Each color and marker type represents a distinct value of the driving time $t_{\rm F}$, as indicated in the bar on the right.
Results of calculations of the target state fidelity $\mathcal{F}_{1}$ from Eq. (24) for drivings across the first-order QPT are summarized in Fig. 3. The two rows of the figure correspond to driving protocols A and B (see Sec. 3.2), columns in each row present results for the three values of the effective system-environment coupling strength $q$ given above. For both $q\neq 0$ values (the weak and strong coupling cases) we distinguish the two choices of the coupling operator $\hat{Q}$ from Eq. (8). Individual curves in each plot correspond to distinct values of the driving time $t_{\rm F}$. These are distinguished by colors and markers that will be used also in some forthcoming figures.
@@
    In the case of protocol A and sufficiently long times $t_{\rm F}$, the onset of interaction with the environment (${q\neq 0}$) leads to a considerable increase of fidelity. The effect is present for both weak and strong coupling, in the latter case being stronger for $\hat{Q}=\hat{J}_{x}$. For protocol B the effect is absent, so the above-mentioned generally better performance of B does not hold for ${q>0}$.
@@
    We observe nonmonotonous dependencies of some of the $\mathcal{F}_{1}$ curves on $T$.
@@
The third instruction is an alternative to the second one and issues from item (c): Stick with driving protocol A but let the system interact with the environment. We see that the slowest driving of type A with the weak coupling ${q>0}$ reaches the fidelity $\mathcal{F}_{1}\approx 0.99$ at low temperature, comparable to driving B with ${q=0}$, but in this case the effect survives even to medium temperatures. The increase of fidelity is present even for the strong S-E coupling, although in a less distinct form and also depending on the coupling operator $\hat{Q}$. The increase of the target ground-state population in these cases is due to the dissipation. It is interesting that the same mechanism does not apparently work for driving protocol B. A possible explanation follows from ESQPT-related avoided crossings of excited states (cf. Fig. 1), which become relevant for ${q\neq 0}$, when there is some energy exchange with the environment, but are not reflected by the variation of the planar speed $u(t)$. Hence the driving protocol B turns out to be more sensitive to the environmental interactions than A.
Figure 4: Evolving occupation probabilities $P_{n}(t)$ of instantaneous energy eigenstates of the ${N=10}$ qubit system for driving protocol A along the path crossing the first-order QPT with ${t_{\rm F}=10^{2.8}}$ and ${T=10^{-0.04}}$. The choices of $q$ and $\hat{Q}$ are indicated. The bottom row depicts the whole low-lying part of the spectrum, with $P_{n}(t)$ encoded in the color of the respective level (see the color bar). The top row shows the occupation probability of the ground state. The vertical dashed lines mark the avoided crossings of the ground and excited states, which are correlated with changes of the occupation probabilities.
@@
For ${q=0}$ only transitions of the Landau-Zener type modify instantaneous populations of individual states, which are initially set to the thermal values ${P_{n}(0)\propto e^{-\beta E_{n}(\mathbb{\Lambda}_{\rm I})}}$. We observe that in this case the ground-state occupation probability $P_{0}(t)$ sharply drops at the QPT-related ground-state avoided level crossing. However, for ${q>0}$, the dissipation and thermalization processes induced by interactions with the environment gradually push the ground-state occupation probability to higher values after the QPT crossing, in some cases going even above the initial population $P_{0}(0)$.
@@
The last and maybe most surprising strategy for increasing the target ground-state fidelity comes from the observation mentioned in the above item (e): Increase the temperature to the value for which $\mathcal{F}_{1}$ is maximal. Indeed, many curves in Fig. 3, particularly those corresponding to smaller driving times, show a clear maximum at a certain optimal temperature ${T_{\rm opt}>0}$. The increase of fidelity reached in this way is not huge (and sometimes no increase is even present), but in some cases it is not negligible, like for driving A with ${t_{\rm F}=10^{2.8}}$ in the strong coupling case with ${\hat{Q}=\hat{J}_{z}}$, when one goes from ${\mathcal{F}_{1}\approx 0.2}$ at low $T$ to ${\mathcal{F}_{1}\approx 0.6}$ at ${T=T_{\rm opt}}$. While for driving protocol A this effect is present for all values of $q$, for protocol B it appears only for ${q>0}$.
Figure 5: Occupation probabilities $P_{n}(t)$ of instantaneous energy eigenstates of the ${N=10}$ qubit system for driving protocol A along the path across the first-order QPT with ${t_{\rm F}=10^{2.8}}$ at temperatures $T$ specified in each plot. The choices of $q$ and $\hat{Q}$ are indicated. All plots in the top row correspond to zero temperature. In the bottom row, the temperature is set to the respective optimal value $T_{\rm opt}$ in the leftmost and righmost plots. In the other cases (for which the dependence of $\mathcal{F}_{1}$ on $T$ in Fig. 3 is monotonous) $T$ is set to the same value as in the rightmost plot.
@@
The nature of this effect is studied in Fig. 5. It compares evolving occupation probabilities (28) of individual energy eigenstates of the qubit system for protocol A with a particular final time $t_{\rm F}$, the driving being performed at different temperatures $T$. We again (as in Fig. 4) observe changes of level populations induced by transitions of the Landau-Zener type and those caused by interactions with the environment. In the top row of plots, which all correspond to ${T=0}$, the depopulation of the ground state in the QPT region via the Landau-Zener mechanism plays an important role for the final fidelity. An interesting exception is the case with ${q=0.1}$ and ${\hat{Q}=\hat{J}_{x}}$, where energy transfers from the environment excite the system even far before the QPT. The bottom row of plots corresponds to the same driving at higher temperatures. We notice that in this case, the thermal population of excited states can sometimes increase the population of the ground state right after the QPT. This happens via the same Landau-Zener mechanism, which now has a partly positive effect. In particular, for ${q=0}$ (the leftmost plot) and for ${q=0.1}$, ${\hat{Q}=\hat{J}_{z}}$ (the rightmost plot) we set $T$ to its respective optimal value $T_{\rm opt}$ and observe an improvement of the final fidelity. However, because of complex interplay between all effects involved in the evolution of individual populations, the increase of temperature does not always lead to an advantage, as seen in Fig. 3. Indeed, for the other choices of $q$ and $\hat{Q}$ in Fig. 5 (the medium three panels in the bottom row), for which $\mathcal{F}_{1}$ in Fig. 3 monotonously decrease with $T$, we detect a lowering of the ground-state population at the end of the driving procedure in comparison with the $T=0$ case.
@@ 4.2 Second-order QPT
Figure 7 displays the target ground-state fidelity $\mathcal{F}_{2}$ from Eq. (3.3) for the path across the second-order QPT. The plots in this figure show all dependencies as in the previous case of the first-order QPT (Sec. 4.1), i.e., the dependence on the temperature $T$, the system-environment coupling strength $q$, coupling type ${\hat{Q}=\hat{J}_{x}}$ and $\hat{J}_{z}$, driving time $t_{\rm F}$, and the driving protocol A and B. Since the minimal ground-state energy gap $\Delta_{10}$ along second-order QPT path decreases with $N$ algebraically, the driving times $t_{\rm F}$ needed to reach values of fidelity comparable to the first-order QPT path (for which the gap closes exponentially) are much shorter than in that case. So we move from the time scale $t_{\rm F}\sim 10^{0.4}\div 10^{3.6}$ used for drivings across the first-order QPT to $t_{\rm F}\sim 10^{0}\div 10^{1.6}$ used for drivings across the second-order QPT.
@@
    The effect of environment-induced increase of fidelity is more or less gone. Interactions with the environment start playing some role only for strong coupling and they mostly decrease the fidelity.
@@
    For the strong coupling, ${q=0.1}$, we observe a strong dependence of $\mathcal{F}_{2}$ on the coupling operator ${\hat{Q}=\hat{J}_{x}}$ and $\hat{J}_{z}$.

### v4

@@ 3.3 Target states and fidelity
The target state for any of the above-described driving procedures is the ground state of the qubit system at the final parameter point ${\mathbb{\Lambda}=\mathbb{\Lambda}_{\rm F}}$. For the path across the first-order QPT with ${\mathbb{\Lambda}_{\rm F}=(0.25,1.2)}$, the final ground state is given by a single eigenvector $|E_{0}(\mathbb{\Lambda}_{\rm F})\rangle$, which is energetically well separated from the other eigenvectors. So, if $\hat{\rho}_{\text{S}}(t_{\rm F})$ is the density operator of the qubit system at the end of the driving procedure, see Eq. (16), then the fidelity of the target-state preparation is
@@
$\mathcal{F}_{2}={\rm tr}\left[\hat{P}_{{\rm gs}}(\mathbb{\Lambda}_{\rm F})\,\hat{\rho}_{\text{S}}(t_{\rm F})\right]=$ (26)
@@ 3.4 Notes on the HEOM calculations
The evolution of the qubit density operator $\hat{\rho}_{\text{S}}(t)$ from Eq. (16) is determined by the HEOM method Tani89 ; Tani06 ; Fruch16 ; Tani20 ; Lamb23 , which is capable to yield exact (nonperturbative) numerical description of the system dynamics. In particular, it goes beyond the approximations involved in the common Lindblad formalism, which strictly relies on the assumption that environment-induced processes on the system are Markovian. The method originates from the influence functional formalism of Feynman and Vernon Feyn63 , which arises from the integration over all degrees of freedom of the environment and determines the system evolution $\hat{\rho}_{\text{S}}(t)$ on the basis of the system self-Hamiltonian $\hat{H}_{\rm S}$, the coupling operator $\hat{Q}$ from Eq. (7) and the environment self-correlation function $C(t)$ from Eq. (11).
We do not immerse here deeper into the description of the HEOM method but just present its very rough outline and introduce the parameters involved. The evolution $\hat{\rho}_{\text{S}}(t)$ is described by a set of coupled first-order differential equations Tani89 , which can be efficiently solved in terms of some auxiliary density matrices. These matrices do not represent real physical states but serve just as a numerical tool to account for the S-E correlations depending on the coupling strength $q$, temperature $T$ and dissipation time scale $\tau_{\rm E}$. When $q$ is small and $\tau_{\rm E}$ much lower than the internal dynamical time scale of the system, the dynamics is approximately Markovian and only a few auxiliary density matrices are needed. In contrast, for large $q$ and $\tau_{\rm E}$ the dynamics is non-Markovian and a large number of matrices is required to account for the long-lasting S-E correlations. The auxiliary matrices are grouped under a certain hierarchy of layers enumerated by ${0,1,\dots,L}$, which are connected with the inclusion of time correlations of an increasing order. The depth $L$ of this hierarchy must be chosen large enough to ensure the convergence. Details can be found, e.g., in Ref. Tani06 .
In order to solve the set of HEOM equations, the self-correlation function (11) is expanded as a series of exponentials. This can be done analytically, in general, for a spectral density $J(\omega)$ that admits a finite number of poles in the complex plane and that goes to zero as $\omega\rightarrow\infty$. For the Drude-Lorentz distribution (13) the series takes the form
@@
with coefficients ${c_{0}\in{\mathbb{C}}}$ and ${c_{1},c_{2},\ldots\in{\mathbb{R}}}$ depending on parameters $q$ and $\gamma$, and on the temperature $T$. The explicit expressions are given, e.g., in Ref. Lamb23 . In numerical calculations the sum in Eq. (27) can be truncated at its $M$th term. For low temperatures, the convergence is slow and $M$ must be relatively large to reach an acceptable precision of calculations. On the other hand, for high temperatures even very small $M$ may lead to reasonable results. In any case, to compensate this truncation, a Lindbladian term that speeds up the convergence (so-called terminator) is included Ishi05 .
@@ 4 RESULTS
In our calculations we use the value ${\gamma=10}$ for the width of the Drude-Lorentz distribution (13) in units of $\varepsilon$. This means that the typical time scale of the environment is smaller (10 times for ${\lambda=\chi=0}$) than that of the qubit system. We consider three values (again in units of $\varepsilon$) of the effective system-environment coupling strengths: ${q=0}$ (no coupling), ${q=0.1}$ (medium coupling) and ${q=1}$ (strong coupling). The above nomenclature concerning the coupling strength should not be taken literally as the effect of the coupling is expected to depend on the driving path. In particular, for the path across the first-order QPT both ${q\neq 0}$ values represent relatively strong coupling regimes (the case ${q=1}$ may even be called an ultrastrong coupling), while for the path across the second-order QPT the case ${q=0.1}$ corresponds to a weak coupling regime. In most calculations, we set the size of the qubit system to a relatively low value, namely ${N=10}$. In that case, the width $\gamma$ of the Drude-Lorentz distribution covers the whole spectrum of the qubit system with ${\lambda=\chi=0}$.
At low temperatures, values of the truncation parameter up to ${M=18}$ are needed to ensure the convergence of the HEOM calculations, while at high temperatures, the value ${M=5}$ is found to be sufficient. For all temperatures, the number of layers is set to ${L=3}$, the contributions of the ${L\geq 4}$ layers being shown to yield only negligible corrections. This holds even at the low temperatures, when a short period of ${{\rm Re}\,C(t)<0}$ induces a strong non-Markovian behavior.
@@
where $(N+1)$ is the dimension of the qubit Hilbert space. Therefore, we consider the fidelity for different driving times $t_{\rm F}$ as a function of temperature ${T\leq 25\approx 10^{1.4}}$.
@@ 4.1 First-order QPT
Figure 3: Final fidelity of the ${N=10}$ qubit system for drivings across the first-order QPT precursor. Results for driving protocols A and B are shown in the top and bottom rows, respectively. The effective coupling $q$ from (10), the angle $\theta$ in (8) and the parameter ${r}$ in (9) are indicated above or within each subplot. Each color and marker type represents a distinct value of the driving time $t_{\rm F}$, as indicated in the bar on the right (the markers are used just to label the curves and do not correspond to all calculated points). The range of temperature is $\log_{10}T\in[-0.8,1.4]$.
Results of calculations of the target state fidelity $\mathcal{F}_{1}$ from Eq. (25) for drivings across the first-order QPT are summarized in Fig. 3. The two rows of the figure correspond to driving protocols A and B (see Sec. 3.2), columns in each row present results for the three values of the effective system-environment coupling strength $q$ given above. For both $q\neq 0$ values (the medium and strong coupling cases) we distinguish the two choices of the coupling operator $\hat{Q}$ from Eq. (8) by the angle $\theta$. The inclusion of the counterterm in (9) is indicated by ${r=1}$ ($r=0$ implies no addition of it). Individual curves in each plot correspond to distinct values of the driving time $t_{\rm F}$. These are distinguished by colors and markers that will be used also in some forthcoming figures.
@@
    In the case of protocol A and sufficiently long times $t_{\rm F}$, the onset of interaction with the environment (${q\neq 0}$) leads to a considerable increase of fidelity. The effect is present for both medium and strong couplings, being more pronounced for ${\theta=\frac{\pi}{2}}$ and ${r=0}$. For protocol B the effect is absent, so the above-mentioned generally better performance of B does not hold for ${q>0}$.
@@
    We observe nonmonotonous dependencies of some of the $\mathcal{F}_{1}$ curves on $T$ for both driving types.
  • (f)
    Effects caused by the inclusion of the counterterm (${r=1}$) in Eq. (9) depend on the coupling strength: For ${q=1}$, we observe a systematic increase of fidelity (compared to the corresponding ${r=0}$ cases) for long times $t_{\rm F}$ and a decrease for shorter times. This holds for both driving types. For ${q=0.1}$ the effect of the counterterm depends nonsystematically on the driving type and time.
@@
Figure 4: Evolving occupation probabilities $P_{n}(t)$ of instantaneous energy eigenstates ($n=0,1,\ldots$) of the ${N=10}$ qubit system for driving protocol A along the path crossing the first-order QPT with ${t_{\rm F}=10^{2.8}}$ and ${T=10^{-0.04}}$. The choices of $q$, $\theta$ and $r$ are indicated. The bottom row depicts the whole low-lying part of the spectrum, with $P_{n}(t)$ encoded in the color of the respective level (see the color bar). The top row shows the occupation probability of the ground state. The vertical dashed lines mark the avoided crossings of the ground and excited states, which are correlated with changes of the occupation probabilities.
The third instruction is an alternative to the second one and issues from item (c): Stick with driving protocol A but let the system interact with the environment. We see that the slowest driving of type A with the S-E coupling strength ${q=0.1}$ reaches the fidelity $\mathcal{F}_{1}\approx 0.99$ at low temperature, comparable to driving B with ${q=0}$, but in this case the effect survives even to medium temperatures. The increase of fidelity is present even for the strong coupling ${q=1}$ , although in a less distinct form and also depending on the coupling operator $\hat{Q}$. For the strong coupling, the counterterm in Eq. (9) decreases the enhancement for fast drivings but increases it for slower drivings. The increase of the target ground-state population in these cases is due to the dissipation and ongoing thermalization of the system during the drive. The same mechanism apparently does not work for the driving protocol B. In this case, the fidelity with ${q=0}$ is already high and the onset of nonzero coupling washes out the advantage of the ${v={\rm const.}}$ strategy because the S-E interaction populates excited states of the system (which makes the ground-state metric less relevant than in the ${q=0}$ case).
@@
So $P_{0}(t)$ is the instantaneous occupation of the ground state, $P_{1}(t)$ is the instantaneous occupation of the first excited state, and so on. For ${q=0}$ only transitions of the Landau-Zener type modify instantaneous populations of individual states, which are initially set to the thermal values ${P_{n}(0)\propto e^{-\beta E_{n}(\mathbb{\Lambda}_{\rm I})}}$. We observe that in this case the ground-state occupation probability $P_{0}(t)$ sharply drops at the QPT-related ground-state avoided level crossing. However, for ${q>0}$, the dissipation and thermalization processes induced by interactions with the environment gradually push the ground-state occupation probability to higher values after the QPT crossing, in some cases going even above the initial population $P_{0}(0)$. This holds for both ${r=0}$ and 1 cases.
@@
The last and maybe most surprising strategy for increasing the target ground-state fidelity comes from the observation mentioned in the above item (e): Increase the temperature to the value for which $\mathcal{F}_{1}$ is maximal. Indeed, many curves in Fig. 3, particularly those corresponding to smaller driving times, show a clear maximum at a certain optimal temperature ${T_{\rm opt}>0}$. The increase of fidelity reached in this way is not huge (and sometimes no increase is even present), but in some cases it is not negligible, like for driving A with ${t_{\rm F}=10^{2.8}}$ in the strong coupling case with ${\theta=0}$, when one goes from ${\mathcal{F}_{1}\approx 0.2}$ at low $T$ to ${\mathcal{F}_{1}\approx 0.6}$ at ${T=T_{\rm opt}}$. While for driving protocol A this effect is present for all values of $q$, for protocol B it appears only for ${q>0}$.
@@
The nature of this effect is studied in Fig. 5. It compares evolving occupation probabilities (29) of individual energy eigenstates of the qubit system for protocol A with a particular final time $t_{\rm F}$, the driving being performed at different temperatures $T$ (this time we only show results without the counterterm). We again (as in Fig. 4) observe changes of level populations induced by transitions of the Landau-Zener type and those caused by interactions with the environment. In the top row of plots, which all correspond to ${T=0}$, the depopulation of the ground state in the QPT region via the Landau-Zener mechanism plays an important role for the final fidelity. An interesting exception is the case with ${q=1}$ and ${\theta=\frac{\pi}{2}}$, where energy transfers from the environment excite the system even far before the QPT. The bottom row of plots corresponds to the same driving at higher temperatures. We notice that in this case, the thermal population of excited states can sometimes increase the population of the ground state right after the QPT. This happens via the same Landau-Zener mechanism, which now has a partly positive effect. This effect is even enhanced by transfers of populations to lower excited states when passing the ESQPT-related sequence of avoided-crossings before the QPT. In particular, for ${q=0}$ and for ${q=1}$, ${\theta=0}$ we set $T$ to its respective optimal value $T_{\rm opt}$ and observe an improvement of the final fidelity. However, because of complex interplay between all effects involved in the evolution of individual populations, the increase of temperature does not always lead to an advantage, as seen in Fig. 3. Indeed, for the other choices of $q$ and $\hat{\theta}$ in Fig. 5, for which $\mathcal{F}_{1}$ in Fig. 3 monotonously decrease with $T$, we detect a lowering of the ground-state population at the end of the driving procedure in comparison with the $T=0$ case.
Figure 5: Occupation probabilities $P_{n}(t)$ of instantaneous low-energy eigenstates of the ${N=10}$ qubit system for driving protocol A along the path across the first-order QPT with ${t_{\rm F}=10^{2.8}}$ at temperatures $T$ specified in each plot. The values of $q$ and $\theta$ are indicated, ${r=0}$. All plots in the top row correspond to zero temperature. In the bottom row, the temperature is set to the respective optimal value $T_{\rm opt}$ for the ${q=0}$ and ${q=1}$, ${\theta=0}$ plots. In the other cases, for which the dependence of $\mathcal{F}_{1}$ on $T$ in Fig. 3 is monotonous, $T$ is set to the same value as in the latter case.
@@ 4.2 Second-order QPT
Figure 7 displays the target ground-state fidelity $\mathcal{F}_{2}$ from Eq. (26) for the path across the second-order QPT. The plots in this figure show all dependencies as in the previous case of the first-order QPT (Sec. 4.1), i.e., the dependence on the temperature $T$, the S-E coupling strength $q$, the angle $\theta$ of the coupling operator (8), the driving time $t_{\rm F}$, the parameter ${r}$ in (9), and on the driving protocols A and B. Since the minimal ground-state energy gap $\Delta_{10}$ along second-order QPT path decreases with $N$ algebraically, the driving times $t_{\rm F}$ needed to reach values of fidelity comparable to the first-order QPT path (for which the gap closes exponentially) are much shorter than in that case. So we move from the time scale ${t_{\rm F}\in[10^{0.4},10^{3.6}]}$ used for drivings across the first-order QPT to ${t_{\rm F}\in[10^{0},10^{1.6}]}$ used for drivings across the second-order QPT.
@@
    The effect of environment-induced increase of fidelity is more or less gone. Interactions with the environment start playing some role only for strong coupling (as expected) and they mostly decrease the fidelity.
@@
    The counterterm in the Hamiltonian improves the fidelity for the strong coupling (compared to the ${r=0}$ case).
  • (g)
    For the strong coupling, ${q=1}$, we observe a strong dependence of $\mathcal{F}_{2}$ on the coupling operator ${\hat{Q}=\hat{J}_{x}}$ (i.e., ${\theta=\frac{\pi}{2}}$) and ${\hat{J}_{z}}$ (i.e., ${\theta=0}$).
The peak of some fidelity curves, see item (e), has been subject to the same analysis as in the first-order QPT case, the result being summarized in Fig. 8. We again observe the linear increase of the optimal temperature $T_{\rm opt}$ and a roughly algebraic decrease of ${\rm max}\,\mathcal{F}_{2}$ with the size $N$ of the qubit system.
In the case of the second-order QPT path, special attention needs to be paid to the choice of the system-environment coupling operator $\hat{Q}$ in Eq. (8). As pointed out in item (g), this choice makes a large difference in the values of $\mathcal{F}_{2}$ obtained for the strong coupling. This effect is due to the difference between both choices in terms of the parity conservation. We know that for ${q=0}$ the driving from the positive-parity initial state along the ${\chi=0}$ path can excite the system only to states with positive parity. However, for ${q\neq 0}$ the parity may be violated by system-environment interaction. This happens for ${\theta=\frac{\pi}{2}}$, so during the drive the environment induces strong transitions to the first excited state with negative parity. This population is then counted in the summed final fidelity $\mathcal{F}_{2}$ in Eq. (26), so for slow-enough driving (when transitions to higher excited states can be neglected) this fidelity may be relatively high.
In contrast, for ${\theta=0}$ the parity is conserved even by the system-environment interaction, so the strongest transitions during the drive lead to the second excited state with positive parity, which does not contribute to $\mathcal{F}_{2}$. This explains why slow drivings in the strong-coupling columns of Fig. 7 yield higher fidelity for ${\theta=\frac{\pi}{2}}$ than for ${\theta=0}$. In the latter case all $\mathcal{F}_{2}$ curves converge to a narrow band, washing out the difference between slow and fast driving. It seems that for ${\theta=0}$ the environment-induced transitions and the transitions of the Landau-Zener type have similar effects which complement each other as the driving time $t_{\rm F}$ varies.
4.3 Comparison with the Lindblad method
In this section we compare the above HEOM results with simulations using the Lindblad method. The use of the Lindblad formalism for time-dependent Hamiltonians is discussed, e.g., in Refs. Alba12 ; Yama17 ; Dann18 . In our calculations we apply the Lindblad formula
$\frac{d}{dt}\hat{\rho}_{\text{S}}(t)=-i[\hat{H}_{\text{S}}(t)+\hat{H}_{\text{L}}(t),\hat{\rho}_{\text{S}}(t)]+\mathcal{D}_{t}\hat{\rho}_{\text{S}}(t)$ (30)
with the dissipator,
$\mathcal{D}_{t}\hat{\rho}_{\text{S}}(t)\!=\!\!\sum_{\epsilon}\!\frac{\Gamma(\epsilon)}{N}\,\!\!\left[\hat{S}(\epsilon)\hat{\rho}_{\text{S}}(t)\hat{S}^{\dagger}\!(\epsilon)\!-\!\frac{1}{2}\big{\{}\hat{S}^{\dagger}\!(\epsilon)\hat{S}(\epsilon),\hat{\rho}_{\text{S}}(t)\big{\}}\right]\!.$ (31)
Here the sum goes over all possible differences $\epsilon={E_{m}(\mathbb{\Lambda}(t))-E_{n}(\mathbb{\Lambda}(t))}$ between instantaneous system energies and
$\hat{S}(\epsilon)$ $=$ $\!\!\!\!\sum\limits_{\begin{subarray}{c}m,n\\ E_{m}-E_{n}=\epsilon\end{subarray}}\!\!\!\!\langle E_{n}|\hat{Q}|E_{m}\rangle|E_{n}\rangle\langle E_{m}|,$ (32)
$\Gamma(\epsilon)$ $=$ $2\big{[}1\!+\!n_{\rm E}(\epsilon)\big{]}\big{[}J(\epsilon)\Theta(\epsilon)\!-\!J(-\epsilon)\Theta(-\epsilon)\big{]},$ (33)
where $\Theta$ denotes the Heaviside step function. The function (33) is positive and satisfies ${\Gamma(0)=4qT/\gamma}$ for the Drude-Lorentz form (13). Negative or positive values of $\epsilon$ indicate excitation or relaxation of the system, respectively, while ${\epsilon=0}$ indicates dephasing Breu02 . The hermitian term
$\hat{H}_{\text{L}}(t)=\frac{1}{N}\!\sum_{\epsilon}\left[\frac{c_{0}\epsilon}{\gamma^{2}+\epsilon^{2}}+\sum_{k=1}^{\infty}\frac{c_{k}\epsilon}{(2\pi kT)^{2}+\epsilon^{2}}\right]\!\hat{S}^{\dagger}\!(\epsilon)\hat{S}(\epsilon)$ (34)
in (30) represents the so-called Lamb shift Hamiltonian, with coefficients $c_{0},c_{1},\ldots$ introduced in (27).